OpenStudy (anonymous):

If the velocity of a particle at time t is represented by the equation v(t) = 8t + 2, find the position of the particle at time 3 if the position at time 0 is 0.

OpenStudy (anonymous):

dx/dt =8t+2 on integrating u get x=4t^2 +2t +c at t=0 x=0 thus we have c=0 thus x=4t^2 +2t at t=3 the position is 4(3)^2+2(3)=42

OpenStudy (anonymous):

is distance(s) and position the same?

OpenStudy (anonymous):

one can assume here

OpenStudy (anonymous):

why did we have to integrate it? and what is x in the problem?

OpenStudy (anonymous):

x is the position or disance in the problem... v=dx/dt so in order to get x we have to integrate v

OpenStudy (anonymous):

ok..thanks a lot!!

OpenStudy (anonymous):

welcome